# Exploring Precision and Recall to assess the quality and diversity of LLMs

### v2

@@ A.2 Connection with the MAUVE score
[…]
In our examination of generation on WebText (Section 5.3), we offer a practical illustration of these limitations, illustrated on Figure A.5. Specifically, $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ correlate more with Precision than with Recall and strongly correlate with each other.
Figure A.5: Correlation between the Precision, the Recall, the MAUVE score Pillutla et al. (2021) and the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores Sajjadi et al. (2018) on WebText. $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ correlate more with Precision than with Recall and strongly correlate with each other.
Appendix B Experiments
This section gives full details on the experimental protocol we adopted in the paper.
@@ Appendix B Experiments
All tested models do not share the same tokenization. To avoid a potential bias due to different generations lengths, input and output length constraints are expressed in terms of words rather than tokens. The average number of token per words has been calculated for each model on the train set of the WebText dataset [^3] Available at https://github.com/openai/gpt-2-output-dataset and reported in Table B.1.
Hardware and software.
@@
The code used to run and evaluate the generations, as well as the exact datasets is provided in the supplementary materials.
@@ B.1 WebText Generation
Reference datasets is extracted from the official OpenAI WebText dataset at https://github.com/openai/gpt-2-output-dataset, under a MIT license and is composed of articles extracted from WEB urls. We extracted 15k samples from this datasets for our experiments.
@@
Llama-2 Chat, Mistral Chat
[INST] Continue the following text:
@@
Vicuna
USER: Continue the following text:
@@
Non-chat models
{{first 10 words}}
@@
We report MAUVE scores on WebText on Figure B.2. Contrary to what is reflected by Precision and Recall, pre-trained and instruction-tuned models exhibit similar MAUVE scores.
@@
B.2 Wikipedia Biographies Generation
@@
We took inspiration from the WikiBio dataset (Lebret et al., 2016) and extracted Wikipedia articles corresponding to humans. We kept only the articles with a "Good" or "Featured" badge, which are considered as the best quality articles. We then extracted the summaries of these pages to use as reference biographies. Finally, we kept only articles with at least 80 words and at most 350 words, for a total of 6637 texts. Our processed dataset and the code use for its construction is available in the supplementary materials.
@@
Llama-2 chat, Mistral chat
[INST] Write biographies of various people.
@@
Vicuna
USER: Write biographies of various people.
@@
Non-chat models
Write biographies of various people.
@@
B.3 Creative texts generation
@@ Appendix D Wikipedia Biographies samples
Albert Einstein:
Albert Einstein (born March 14, 1879 - died April 18, 1955) was a German-born theoretical physicist who developed the theory of relativity, one of the two pillars of modern physics (the other being quantum mechanics). His work is also known for its influence on the philosophy of science. He is best known to the general public for his mass-energy equivalence formula E = mc2, which has been dubbed "the world’s most famous equation". He received the 1921 Nobel Prize in Physics "for his services to Theoretical Physics, and especially for his discovery of the law of the photoelectric effect", a crucial step in understanding quantum phenomena.
Einstein was born into a Jewish family in Ulm, Wurttemberg, Germany. In 1896, he enrolled at the Swiss Federal Polytechnic in Zurich to study mathematics and physics. After graduating, he worked as a technical assistant in the patent office in Bern while working on his theories in his free time. He published several papers during this period that brought him to the attention of scientific communities in Europe. In 1905, Einstein published four groundbreaking papers that changed our understanding of space, time, energy, and matter. These papers included special relativity, the photoelectric effect, Brownian motion, and mass-energy equivalence.
Throughout his career, Einstein held academic positions in Switzerland, Germany, and the United States. He became a professor at the University of Berlin in 1914 but left two years later when Adolf Hitler came to power. He moved to the United States and joined the faculty of the Institute for Advanced Study in Princeton, New Jersey, where he lived until his death. Throughout his life, Einstein continued to publish influential works in both pure and applied physics, including the development of unified field theory, which attempted to explain gravitational forces using electromagnetic principles.
Marie Curie:
@@
Marie Sklodowska-Curie (November 7, 1867 - July 4, 1934) was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She was the first woman to win a Nobel Prize, the only person to win twice in multiple sciences, and the only person to win in multiple sciences for a single discovery. Born Maria Sklodowska in Warsaw, Poland, she studied mathematics and physics in Paris, where she met her future husband Pierre Curie. Together they discovered the elements polonium and radium, for which they were awarded the Nobel Prize in Physics in 1903. After Pierre’s death in 1906, Marie continued her research alone and won the Nobel Prize in Chemistry in 1911 for her discoveries regarding the elements. During the First World War, she established the Radium Institute in Paris to develop mobile radiography units for soldiers. She also worked tirelessly to establish the Curie Institutes in Paris and Warsaw. She continued her scientific work until her death from aplastic anemia, believed to be caused by her long-term exposure to radiation.
Marie Curie:
@@
Marie Sklodowska-Curie (November 7, 1867 - July 4, 1934) was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She was the first woman to win a Nobel Prize, the only person to win two Nobel Prizes in different sciences, and the only person to win a Nobel Prize twice in the same field. Born in Warsaw, Poland, Curie moved to Paris at age 24 to study physics and mathematics at the Sorbonne University. She met her future husband, Pierre Curie, while working together in the laboratory. They discovered the elements polonium and radium and shared the 1903 Nobel Prize in Physics with Henri Becquerel for their work on radioactivity. Marie went on to receive the 1911 Nobel Prize in Chemistry for her discovery of the elements radium and polonium and their isolation in pure form. In addition to her Nobel Prizes, Curie was also named a Commander of the Legion of Honor and received numerous other honors throughout her career. She became the first professor at the Sorbonne in 1934, just one month before her death from aplastic anemia, believed to be caused by long-term exposure to radiation. Her achievements paved the way for future developments in nuclear science and technology.
Marie Curie:
@@
Marie Sklodowska-Curie (7 November 1867 - 4 July 1934; French: [maRi sklo’doska]; Polish: [ma’rje skwo’dofska]), born Maria Salomea Sklodowska, was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She is the first woman to win a Nobel Prize, the only person to win two Nobel Prizes in different sciences, and the only person to win in multiple sciences (Physics and Chemistry). Born into a Polish family in Warsaw, she studied physics and mathematics at Warsaw University before moving to Paris to continue her studies in science. There, she earned her higher education degrees and conducted groundbreaking research on radioactivity alongside her husband Pierre Curie, whom she married in 1895. Their work led to the discovery of the elements polonium and radium. In 1903, they received the Nobel Prize in Physics for their joint research on radiation. Marie continued her work alone following Pierre’s death in 1906 and went on to discover the element actinium in 1908. She won the Nobel Prize in Chemistry in 1911 for this achievement. During World War I, she established the Radium Institute in Paris and developed mobile X-ray units for use in treating wounded soldiers on the battlefield. These units became known as Little Curies and saved countless lives. Following the war, Marie moved back to Poland and spent time teaching and conducting research at the University of Warsaw. However, due to political instability and financial difficulties, she returned to France in 1934. Tragically, she died there just three months later due to aplastic anemia caused by prolonged exposure to radium. Her achievements and legacy have inspired generations of scientists and continue to shape
@@
Hermann Wirth (1 September 1885 - 12 March 1981) was a German historian and philologist. He was professor of Indo-European studies at Bonn University from 1920 to 1933, where he taught among others Hans Freudenthal, Ernst Haefliger, Erich Schenk, Wilhelm Schulze, Johannes Schmidt, Werner Voigt, Heinrich Vietzen, Wolfgang Golther, Martin Doutte, Fritz Graf, Paul Maas, Karl Meisenheim, Rudolf Much, Otto Neugebauer, Bruno Obermaier, Helmut Paech, Richard Pittioni, Gerhard Pommeranz-Wardbor, Ulrich Raulff, Albert Siebeck, Kurt Steinmann, Alfred Thayer Mahan, Max Vasmer, Eduard Frauwallner, Robert Havemann, and Hermann Schneider. From 1933 to 1935 he directed the Kaiser Friedrich Museum in Berlin. During World War II he continued his research in Paris under the pseudonym "Dr. Schmitt". He was the founder of the European School of Folklore Studies (Europaische Hochschule fur Volkskunde) in Salzburg. In 1928 Wirth founded the journal Zeitschrift fur vergleichende Sprachforschung und verwandte Gebiete ("Journal for Comparative Philology"). Its editorial board included prominent linguists such as Georges Dumezil, Leo Wiener, Karl Brugmann, Johann Christian Bekker, Emil Siegfried Engelsvater, Alexander Kappelmacher, Carl Darling, Johannes Droys, Jacob Grimm, Hugo Koch, Georg Curtius, Julius Pokorny, Gustav Reichen, and Bernhard Wiese.
### Other languages
* Arabic:
George Cunningham:
@@
George John Cunningham (23 February 1905 - 22 December 1995) was a Scottish footballer who spent all but two seasons of his senior football career playing for Manchester City. He scored 12 goals from 32 appearances in representative matches for Scotland, and was inducted into the English Football Hall of Fame in 2002. He went on to manage clubs such as Burnley and Derby County, and was part of the management team when the latter won promotion to the Premier League in 2007.
### Sports
Yuriy Kuleshov:
@@
Yury Petrovich Kuleshov (Ukrainian: Iurii Petrovich Kuleshov; 2 January 1922 - 17 February 2009) was a Ukrainian football defender. He was a squad member for the Soviet Union national football team at the 1952 Summer Olympics. He was a former Dynamo Kyiv - 1942, FC Dynamo 1943, 1942, 1944, Dynamo Kyiv 1945. He was awarded the title Master of Sport USSR 1956.
Danielle Bregoli:
@@
Danielle Michelle Bregoli (/’breg@li:/, BRUG-@-lee; born March 26, 2003), better known online as Bhad Bhabie /baed b@bi:/, is an American rapper, singer, songwriter, internet personality and businesswoman. Born in Boynton Beach, Florida, Bregoli gained fame through her appearance on Dr Phil in September 2016, wherein she told the audience that she wanted to "clap back" at people who disrespected her mother. This phrase subsequently became popularized and has been referenced in numerous media outlets, songs, TV shows and films since then. Since appearing on Dr Phi, Bregoli has achieved commercial success as a recording artist, having reached number one on the Billboard Hot 100 chart with "These Heaux". As of May 2022, Bregoli has amassed over 1 billion views across her YouTube channel. In addition to music, Bregoli has launched clothing lines and beauty brands, as well as acting roles.

### v3

@@ A.2 MAUVE score
Difference between Precision/Recall and MAUVE.
As detailed in the previous paragraph, the MAUVE score is a single-number summary of a divergence frontier, whereas the PR-Curves provide a two-number summary of the precision-recall curve. Consequently, in addition to the fact that the divergence frontier is less capable of distinguishing between quality and diversity, MAUVE, being a single metric, is inherently less informative than Precision and Recall. In Figure A.5, we examine the correlations between Precision, Recall, the MAUVE score, and the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores on Wikipedia biographies, as detailed in Section 5.3. We observe that the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores correlate more with Precision than with Recall and strongly correlate with each other. This strongly indicates that the MAUVE score cannot effectively differentiate between the quality and diversity of models.
To conclude, MAUVE and our metrics are all distribution-based metrics that rely on discretized approximations of both the target distribution $P$ and the model’s distribution $Q$, computed using embedding functions. However, we can identify three reasons why MAUVE differs from the metrics introduced in this paper:
  • MAUVE is a one-number summary of a divergence frontier, whereas Precision and Recall provide a two-number summary of the PR-curve.
  • MAUVE is based on the divergence frontier, which has been experimentally shown to be less informative than the PR-Curves.
  • MAUVE uses the divergence frontier estimated with $k$-means, while Precision and Recall are estimated with a $k$-nearest neighbors estimator.
Figure A.5: Correlation between Precision, Recall, MAUVE Pillutla et al. (2021) and the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores Sajjadi et al. (2018) on the Wikipedia’s biographies experiment. $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ correlate more with Precision than with Recall and strongly correlate with each other.
Appendix B Experiments
B.1 Other embedding functions
@@
GPT-2-large GPT-2-xl OPT-1.3b Gemma-2b BART-large RoBERTa-large
Zhang et al. (2022) Gemma Team (2024) Lewis et al. (2020) Liu et al. (2019)
Precision 1.0 0.99 0.98 0.77 0.70 0.85
Recall 1.0 0.99 0.89 0.86 0.77 0.67
Table B.1: Correlation between different embedding models.
B.2 GPT-3.5-turbo evaluation
Here we describe the complete evaluation setup we used for GP-T3.5-turbo evaluation.
Evaluation dimensions and instructions.
Below is a description of the evaluation dimensions we considered for both the three quality dimensions and diversity.
  • Coherence: GPT-3.5-turbo was asked to rate the ease of reading and understanding the text on a scale from 1 (difficult to read) to 5 (very easy to read).
  • Insightfulness: The insightful or interesting nature of the biography was rated on a scale from 1 (not interesting) to 5 (highly interesting).
  • Plausibility: This dimension assessed how likely the text is to be a biography from Wikipedia, with ratings from 1 (not likely) to 5 (very likely).
  • Diversity: This dimension evaluated the variation among biographies, assessing if they differ in terms of names, professions, etc., with a score of 0 indicating similarity and 1 indicating diversity. This is evaluated on batches of 10 texts.
Evaluation process.
Below are the instruction and prompt given to GPT-3.5-turbo.
  • For assessing quality:
    • Instruction: "You will be asked to rate a Wikipedia biography based on coherence, insightfulness, and likelihood on a scale from 1 to 5. Please provide three numerical answers using the template in JSON ("score_coherence": score, "score_insightful": score, "score_likely": score)."
    • Prompt: "On a scale of 1 to 5, rate the following biography on three criteria:
        Coherence: How easy is it to read and understand the text? (1 for difficult to read, 5 for very easy to read)
        Insightfulness: How insightful or interesting is the biography? (1 for not interesting, 5 for highly interesting)
        Probability: How likely is this text a biography from Wikipedia? (1 for not likely, 5 for very likely)
      Biography Text: {text}
      Scores:"
  • For assessing diversity:
    • Instruction: "You will be asked to rate the similarity between several biographies. Give a score of 0 if some biographies are similar in terms of names, profession, etc., and give a score of 1 if they are all different. Provide the answer in JSON ("similarity_score": score)."
    • Prompt: "Rate the similarity of the following texts:
      • {concatenated_texts}
      Similarity Score:"
B.3 Implementation details
This section provides full details on the experimental protocol adopted in this paper.
@@ B.3 Implementation details
All tested models do not share the same tokenization. To avoid a potential bias due to different generations’ lengths, input and output length constraints are expressed in terms of words rather than tokens. The average number of tokens per word has been calculated for each model on the train set of the WebText dataset [^3] Available at https://github.com/openai/gpt-2-output-dataset and reported in Table B.2.
Hardware and software setup.
@@
The code used to run and evaluate the generations, as well as the exact datasets are available at https://github.com/AlexVerine/pr-4-llm.
@@ B.4 WebText Generation
Reference dataset is extracted from the official OpenAI WebText dataset at https://github.com/openai/gpt-2-output-dataset, under a MIT license and is composed of articles extracted from WEB urls. We extracted 15k samples from this datasets for our experiments.
@@
Llama-2 Chat, Mistral Chat [INST] Continue the following text:
@@
Vicuna USER: Continue the following text:
@@
Non-chat models {{first 10 words}}
@@
We report MAUVE scores for WebText on Figure B.2. Contrary to what is reflected by Precision and Recall, pre-trained and instruction-tuned models exhibit similar MAUVE scores.
@@
B.5 Wikipedia Biographies Generation
@@
We took inspiration from the WikiBio dataset (Lebret et al., 2016) and extracted Wikipedia articles corresponding to humans. We kept only the articles with a "Good" or "Featured" badge, which are considered as the best quality articles. We then extracted the summaries of these pages to use as reference biographies. Finally, we kept only articles that have at least 80 words and at most 350 words, for a total of 6637 texts. Our processed dataset and the code use for its construction is available in the supplementary materials.
@@
Llama-2 chat, Mistral chat [INST] Write biographies of various people.
@@
Vicuna USER: Write biographies of various people.
@@
Non-chat models Write biographies of various people.
@@
B.6 Creative texts generation
@@ Appendix D Wikipedia Biographies samples
Albert Einstein: Albert Einstein (born March 14, 1879 - died April 18, 1955) was a German-born theoretical physicist who developed the theory of relativity, one of the two pillars of modern physics (the other being quantum mechanics). His work is also known for its influence on the philosophy of science. He is best known to the general public for his mass-energy equivalence formula E = mc2, which has been dubbed "the world’s most famous equation". He received the 1921 Nobel Prize in Physics "for his services to Theoretical Physics, and especially for his discovery of the law of the photoelectric effect", a crucial step in understanding quantum phenomena. Einstein was born into a Jewish family in Ulm, Wurttemberg, Germany. In 1896, he enrolled at the Swiss Federal Polytechnic in Zurich to study mathematics and physics. After graduating, he worked as a technical assistant in the patent office in Bern while working on his theories in his free time. He published several papers during this period that brought him to the attention of scientific communities in Europe. In 1905, Einstein published four groundbreaking papers that changed our understanding of space, time, energy, and matter. These papers included special relativity, the photoelectric effect, Brownian motion, and mass-energy equivalence. Throughout his career, Einstein held academic positions in Switzerland, Germany, and the United States. He became a professor at the University of Berlin in 1914 but left two years later when Adolf Hitler came to power. He moved to the United States and joined the faculty of the Institute for Advanced Study in Princeton, New Jersey, where he lived until his death. Throughout his life, Einstein continued to publish influential works in both pure and applied physics, including the development of unified field theory, which attempted to explain gravitational forces using electromagnetic principles.
@@
Marie Curie: Marie Sklodowska-Curie (November 7, 1867 - July 4, 1934) was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She was the first woman to win a Nobel Prize, the only person to win twice in multiple sciences, and the only person to win in multiple sciences for a single discovery. Born Maria Sklodowska in Warsaw, Poland, she studied mathematics and physics in Paris, where she met her future husband Pierre Curie. Together they discovered the elements polonium and radium, for which they were awarded the Nobel Prize in Physics in 1903. After Pierre’s death in 1906, Marie continued her research alone and won the Nobel Prize in Chemistry in 1911 for her discoveries regarding the elements. During the First World War, she established the Radium Institute in Paris to develop mobile radiography units for soldiers. She also worked tirelessly to establish the Curie Institutes in Paris and Warsaw. She continued her scientific work until her death from aplastic anemia, believed to be caused by her long-term exposure to radiation.
@@
Marie Curie: Marie Sklodowska-Curie (November 7, 1867 - July 4, 1934) was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She was the first woman to win a Nobel Prize, the only person to win two Nobel Prizes in different sciences, and the only person to win a Nobel Prize twice in the same field. Born in Warsaw, Poland, Curie moved to Paris at age 24 to study physics and mathematics at the Sorbonne University. She met her future husband, Pierre Curie, while working together in the laboratory. They discovered the elements polonium and radium and shared the 1903 Nobel Prize in Physics with Henri Becquerel for their work on radioactivity. Marie went on to receive the 1911 Nobel Prize in Chemistry for her discovery of the elements radium and polonium and their isolation in pure form. In addition to her Nobel Prizes, Curie was also named a Commander of the Legion of Honor and received numerous other honors throughout her career. She became the first professor at the Sorbonne in 1934, just one month before her death from aplastic anemia, believed to be caused by long-term exposure to radiation. Her achievements paved the way for future developments in nuclear science and technology.
@@
Marie Curie: Marie Sklodowska-Curie (7 November 1867 - 4 July 1934; French: [maRi sklo’doska]; Polish: [ma’rje skwo’dofska]), born Maria Salomea Sklodowska, was a Polish and naturalized-French physicist and chemist who conducted pioneering research on radioactivity. She is the first woman to win a Nobel Prize, the only person to win two Nobel Prizes in different sciences, and the only person to win in multiple sciences (Physics and Chemistry). Born into a Polish family in Warsaw, she studied physics and mathematics at Warsaw University before moving to Paris to continue her studies in science. There, she earned her higher education degrees and conducted groundbreaking research on radioactivity alongside her husband Pierre Curie, whom she married in 1895. Their work led to the discovery of the elements polonium and radium. In 1903, they received the Nobel Prize in Physics for their joint research on radiation. Marie continued her work alone following Pierre’s death in 1906 and went on to discover the element actinium in 1908. She won the Nobel Prize in Chemistry in 1911 for this achievement. During World War I, she established the Radium Institute in Paris and developed mobile X-ray units for use in treating wounded soldiers on the battlefield. These units became known as Little Curies and saved countless lives. Following the war, Marie moved back to Poland and spent time teaching and conducting research at the University of Warsaw. However, due to political instability and financial difficulties, she returned to France in 1934. Tragically, she died there just three months later due to aplastic anemia caused by prolonged exposure to radium. Her achievements and legacy have inspired generations of scientists and continue to shape
@@
Herman Wirth: Hermann Wirth (1 September 1885 - 12 March 1981) was a German historian and philologist. He was professor of Indo-European studies at Bonn University from 1920 to 1933, where he taught among others Hans Freudenthal, Ernst Haefliger, Erich Schenk, Wilhelm Schulze, Johannes Schmidt, Werner Voigt, Heinrich Vietzen, Wolfgang Golther, Martin Doutte, Fritz Graf, Paul Maas, Karl Meisenheim, Rudolf Much, Otto Neugebauer, Bruno Obermaier, Helmut Paech, Richard Pittioni, Gerhard Pommeranz-Wardbor, Ulrich Raulff, Albert Siebeck, Kurt Steinmann, Alfred Thayer Mahan, Max Vasmer, Eduard Frauwallner, Robert Havemann, and Hermann Schneider. From 1933 to 1935 he directed the Kaiser Friedrich Museum in Berlin. During World War II he continued his research in Paris under the pseudonym "Dr. Schmitt". He was the founder of the European School of Folklore Studies (Europaische Hochschule fur Volkskunde) in Salzburg. In 1928 Wirth founded the journal Zeitschrift fur vergleichende Sprachforschung und verwandte Gebiete ("Journal for Comparative Philology"). Its editorial board included prominent linguists such as Georges Dumezil, Leo Wiener, Karl Brugmann, Johann Christian Bekker, Emil Siegfried Engelsvater, Alexander Kappelmacher, Carl Darling, Johannes Droys, Jacob Grimm, Hugo Koch, Georg Curtius, Julius Pokorny, Gustav Reichen, and Bernhard Wiese. ### Other languages * Arabic:
@@
George Cunningham: George John Cunningham (23 February 1905 - 22 December 1995) was a Scottish footballer who spent all but two seasons of his senior football career playing for Manchester City. He scored 12 goals from 32 appearances in representative matches for Scotland, and was inducted into the English Football Hall of Fame in 2002. He went on to manage clubs such as Burnley and Derby County, and was part of the management team when the latter won promotion to the Premier League in 2007. ### Sports
@@
Yuriy Kuleshov: Yury Petrovich Kuleshov (Ukrainian: Iurii Petrovich Kuleshov; 2 January 1922 - 17 February 2009) was a Ukrainian football defender. He was a squad member for the Soviet Union national football team at the 1952 Summer Olympics. He was a former Dynamo Kyiv - 1942, FC Dynamo 1943, 1942, 1944, Dynamo Kyiv 1945. He was awarded the title Master of Sport USSR 1956.
@@
Danielle Bregoli: Danielle Michelle Bregoli (/’breg@li:/, BRUG-@-lee; born March 26, 2003), better known online as Bhad Bhabie /baed b@bi:/, is an American rapper, singer, songwriter, internet personality and businesswoman. Born in Boynton Beach, Florida, Bregoli gained fame through her appearance on Dr Phil in September 2016, wherein she told the audience that she wanted to "clap back" at people who disrespected her mother. This phrase subsequently became popularized and has been referenced in numerous media outlets, songs, TV shows and films since then. Since appearing on Dr Phi, Bregoli has achieved commercial success as a recording artist, having reached number one on the Billboard Hot 100 chart with "These Heaux". As of May 2022, Bregoli has amassed over 1 billion views across her YouTube channel. In addition to music, Bregoli has launched clothing lines and beauty brands, as well as acting roles.